# HOLISMOKES XVI: Lens search in HSC-PDR3 with a neural network committee and post-processing for false-positive removal

### v1

@@
HOLISMOKES XVI: Lens search in HSC-PDR3 with a neural network committee and post-processing for false-positive removal
We have carried out a systematic search for galaxy-scale lenses exploiting multi-band imaging data from the third public data release of the Hyper Suprime-Cam (HSC) survey with the focus on false-positive removal, after applying deep learning classifiers to all 110 million sources with -Kron radius above 08. To improve the performance, we tested the combination of multiple networks from our previous lens search projects and found the best performance by averaging the scores from five of our networks. Although this ensemble network leads already to a false-positive rate (FPR) of 0.01% at a true-positive rate (TPR) of 75% on known real lenses, we have elaborated techniques to further clean the network candidate list before visual inspection. In detail, we tested the rejection using SExtractor and the modeling network from HOLISMOKES IX, which resulted together in a candidate rejection of 29% without lowering the TPR. After the initial visual inspection stage to remove obvious non-lenses, 3,408 lens candidates of the 110 million parent sample remained. We carried out a comprehensive multi-stage visual inspection involving eight individuals and identified finally 95 grade A (average grade ) and 503 grade B () lens candidates, including 92 discoveries showing clear lensing features that are reported for the first time. This inspection also incorporated a novel environmental characterization using histograms of photometric redshifts. We publicly release the average grades, mass model predictions, and environment characterization of all visually inspected candidates, while including references for previously discovered systems, which makes this catalog one of the largest compilation of known lenses. The results demonstrate that (1) the combination of multiple networks enhances the selection performance and (2) both automated masking tools as well as modeling networks, which can be easily applied to hundreds of thousands of network candidates expected in the near future of wide-field imaging surveys, help reduce the number of false positives that is the main limitation in lens search to date.
@@
Strong gravitational lensing emerged in the past decades as a powerful tool to probe galaxy evolution and cosmology. It allows us to obtain in a very precise way the total mass (i.e., baryonic and dark matter) of the galaxy or galaxy cluster acting as the lens (e.g., Bolton et al., 2008; Shu et al., 2017; Caminha et al., 2019) and, by assuming that mass follows light, we can disentangle the mass components and obtain unique insights into the dark matter (DM) distribution (e.g., Schuldt et al., 2019; Shajib et al., 2021; Wang et al., 2022) or DM substructure (e.g., Enzi et al., 2024; Ertl et al., 2024; Lange et al., 2024). Thanks to the lensing magnification, strong lensing also allows us to study high-redshift sources not visible otherwise (e.g., Shu et al., 2018; Vanzella et al., 2021; Meštrić et al., 2022; Stiavelli et al., 2023; Morishita et al., 2024).
@@
In case of a time variable background object, such as a supernova (SN) or a quasar, time delays can be measured (e.g., Courbin et al., 2018; Millon et al., 2020) and exploited for competitive measurements of the value of the Hubble constant (Refsdal, 1964). Given the rarity of SNe and strong lensing events, this time-delay cosmography (TDC) technique was mostly carried out with quasars (e.g., Wong et al., 2020; Acebron et al., 2022; Shajib et al., 2022; Acebron et al., 2024). To date only three strongly lensed SNe are known with time delays usable for a precise measurement of : SN Refsdal lensed by the cluster MACS J1149.52223 (e.g., Grillo et al., 2018; Kelly et al., 2023; Grillo et al., 2024), SN H0pe strongly lensed by the cluster PLCK G165.767.0 (Pascale et al., 2025), and SN Encore with SN Requiem both lensed by the same cluster MACS J01382155 (Ertl et al. in prep., Suyu et al. in prep., Pierel et al. in prep). To prepare a systematic search with the current and upcoming wide-field imaging surveys, we initiated the Highly Optimized Lensing Investigations of Supernovae, Microlensing Objects, and Kinematics of Ellipticals and Spirals (HOLISMOKES Suyu et al., 2020) program. As a precursor for the Legacy Survey of Space and Time (LSST) of the Vera C. Rubin Observatory (Ivezic et al., 2008; Ivezić et al., 2019), we are currently exploiting data from the Hyper Suprime-Cam (HSC), which are expected to be very similar.
@@
In Cañameras et al. (2021, hereafter C21) we presented a residual neural network to search broadly for any static lens, and complemented this in Shu et al. (2022, hereafter S22) with lenses/deflectors at relatively high redshift () to better cover the whole redshift range. Both projects relied on a single convolutional neural network (CNN) and targeted HSC images of the second public data release (PDR2). In this work, we combine multiple networks into an ensemble network, relax our restrictions on filter coverage to increase the observed sky area (requesting bands instead of ), and consider data from PDR3 with slightly deeper and larger footprint.
@@
While the resulting sample of network candidates from HSC is small enough for visual inspection, this will not be the case for LSST, Euclid (Laureijs et al., 2011), Roman (Spergel et al., 2015) and the Chinese Space Station Telescope (Gong et al., 2019), delivering more than a billion images, even with a citizen science approach (e.g., Holloway et al., 2024, Holloway et al in prep., Walmsley et al. in prep.). Consequently, we urgently need further automated ways to lower the false-positive rate (FPR) before an unavoidable visual inspection.
@@
In this paper, we show our deep learning ensemble network and, following C21, apply it to cutouts of any object with -Kron radius above 08 that pass standardised HSC image quality flags (see C21 for details). We explore two different approaches to reject false positives: (1) we run SExtractor to reject images with artefacts or without a real astrophysical source in the cutout, followed by a exclusion through HSC-pixel flags, and (2), we run the residual neural network from Schuldt et al. (2023a, hereafter S23a) to reject systems based on the mass model predictions. Specifically, Sect. 2 presents the tested network committee and their performances on known real lenses. We then describe the two approaches of contaminant rejection in Sect. 3, before we carry out a visual inspection as described in Sect. 4 and present the newly discovered lens candidates. Finally, we give a summary and conclusion in Sect. 5. We exploit trained networks from C21, S22, and Cañameras et al. (2024, hereafter C24), and adopt consequently also a flat CDM cosmology with = 0.32 (Planck Collaboration et al., 2020) and (Bonvin et al., 2017).
@@ 2 Inference with network committees
We followed the approach described in C21 to classify the 110 million galaxies from HSC Wide PDR3 with -band Kron radius 08 and optimize the contamination rate of the candidate strong lens sample. Based on test sets drawn from HSC Wide PDR2 images and designed to closely match a real classification setup, Cañameras et al. (2024) show that the purity and overall classification performance are significantly improved with committees of multiple neural networks (see also e.g., Andika et al., 2023). The highest gain is obtained when combining networks trained on different ground-truth data sets, with different prescriptions for the parameters’ distributions over the mock lenses used as positive examples. Taking the average or multiplication of output scores from networks having little overlap in false positives due to their internal representations is best for improving the true-positive rate (TPR) at low FPR.
@@
We investigated several combinations of neural networks chosen among the best performing ResNet and classical CNNs from C21, S22, and C24. After classifying the 110 million HSC PDR3 image cutouts with these networks, we compared the TPR over three independent sets of confirmed and candidate strong lenses as a function of the total number of lens candidates predicted by the committee. This allowed us to find the optimal committee that maximizes the TPR at a given output sample size (i.e. corresponding to a fixed human inspection time). The first set of test lenses (set 1) includes the 1,249 galaxy-scale grade A or B candidates from the SuGOHI papers 111see https://www-utap.phys.s.u-tokyo.ac.jp/ oguri/sugohi/. The second set (set 2) focuses on 201 galaxy-scale grade A from SuGOHI, which are also included in set 1. The third set (set 3) corresponds to the cleaned set of 178 galaxy-scale grade A or B lens candidates used in C21 C24, which excludes visually identified systems with multiple lens galaxies, or significant perturbation from the environment.
@@
We investigated various combinations of different networks by taking the average or the product of their individual network scores. As examples, the performance of combinations with two, three, four, and five networks are shown in Fig. 1. The best committee that maximizes the TPR in all three test sets at fixed number of candidates was obtained by averaging the scores of five individual networks: the ResNet from C21, Classifier-1 from S22, and three additional ResNets from C24, as illustrated in Fig. 2. We adopted a threshold on the average score of 0.55, which corresponds to 22,393 strong-lens candidates, and a TPR of 51%, 54%, and 75% over test set 1, 2, and 3, respectively. Since the TPR curves reach a plateau for all of the three considered test sets, decreasing the score threshold to include 50,000 additional candidates would have improved the TPR by only 4–5%, which does not justify the substantial increase in human inspection time. Having various test sets was important to check that the plateau in TPR is reached irrespective of the exact set-up of the test set. Since set 3 is representing best the galaxy-galaxy scale lenses we target with this network committee (e.g., set 3 is cleaned from group and cluster lenses), the network committee shows the best performance on this set as expected. Finding about 20,000 candidates among 110 million sources from the parent sample is consistent with the FPR in the range 0.01–0.03% predicted in C24 for networks committees, at a TPR of 75% evaluated among set 3.
@@
The five networks are all from the same type of architecture and include residual neural network blocks (He et al., 2015). Specifically, the Classifier-1 from S22 is based on the CMU DeepLens package from Lanusse et al. (2018), while the other four networks are adapted from the ResNet-18 architecture. The networks were however trained on different ground-truth data sets. We introduce the general properties of the training sets and refer the reader to C21, S22, and C24 for further details.
@@
For all five networks, the realistic mock lenses used as positive examples were simulated with the pipeline described in Schuldt et al. (2021, 2023b). Briefly, the pipeline paints lensed arcs on HSC images of massive luminous red galaxies (LRGs) using Singular Isothermal Ellipsoid (SIE) lens mass profiles, and SIE parameter values inferred from SDSS spectroscopic redshifts and velocity-dispersion measurements. After the high-redshift background sources are drawn from the Hubble Ultra-Deep Field (HUDF, Inami et al., 2017), mock lensed arcs are computed with Glee (Suyu & Halkola, 2010; Suyu et al., 2012b), and convolved with the HSC point spread function, before coaddition with the lens galaxy cutout. Positive examples used to train the ResNet from C21 and Classifier-1 from S22 were drawn from the same parent set of mocks, with (i) nearly-uniform Einstein radius distribution between 075 and 25, (ii) a boosted fraction of lens galaxies at with respect to the parent SDSS sample, (iii) lensed images having , and (iv) a minimal ratio between the brightest arc pixel and the local sky background over the lens LRG cutout. The other three high-performing networks from C24 that are part of the committee were trained on mocks produced with a similar procedure, but with (i) a boosted fraction of red HUDF sources, (ii) a natural Einstein radius distribution between 075 and 25 instead of uniform distribution, and (iii) no lower limit on and a balanced fraction of double and quad configurations. This corresponds to sets labeled L1, L4, and L6 in C24.
@@
In terms of negative examples, four out of the five networks include a mix of 33% spirals, 27% isolated LRGs without arcs, 6% groups, and 33% random galaxies over the HSC footprint, as defined for set N1 in C24. The fifth network, namely Classifier-1 from S22, was primarily targeting high-redshift strong-lenses and trained on negative examples drawn from a parent sample with red () and () colors. All five networks were trained and validated on images in bands, such that we require only the availability of bands in HSC while the image in or can be missing.
@@ 3 Cleaning the output candidate list
Before conducting a visual inspection (see Sect. 4) to identify the high-quality strong lens candidates, the catalog of 22,393 candidates was post-processed to remove obvious artefacts and non-lenses. Here we applied mainly two criteria, as detailed below. An overview table of the stages and resulting sample sizes is given in Table 1.
@@ 3.1 Cleaning using SExtractor and HSC pixel level flags
A substantial fraction of contaminants correspond to cutouts with residual background emission in one or all bands (see some examples in the top row of Fig. 3). These cutouts were identified using source masking and estimates of the sky background with SExtractor (Bertin & Arnouts, 1996), leaving 19,820 candidates without loss of any test lens. This step was then refined with information from the pixel-level flags inferred by the HSC pipeline. We searched for optimal cuts to remove contaminants using the flags in bands over the 72 72 pixel images. First, we discarded cutouts having 90% pixels with flags 512 and 1024, corresponding to pixels located near a bright object. This securely excluded cutouts affected by bright neighboring stars within a few tens of arcseconds. Secondly, we removed cutouts with 99% pixels flagged as “detected pixels”, which are more likely to correspond to crowded fields such as stellar clusters, or star-forming clumps within extended disks, than strong gravitational lenses. These two criteria decreased the sample to 18,712 candidates.

### v2

# HOLISMOKES XVI. Lens search in HSC-PDR3 with a neural network committee and post-processing for false-positive removal

@@
HOLISMOKES XVI. Lens search in HSC-PDR3 with a neural network committee and post-processing for false-positive removal
We have carried out a systematic search for galaxy-scale lenses exploiting multiband imaging data from the third public data release of the Hyper Suprime-Cam (HSC) survey with the focus on false-positive removal, after applying deep learning classifiers to all 110 million sources with an -Kron radius above 08. To improve the performance, we tested the combination of multiple networks from our previous lens search projects and found the best performance by averaging the scores from five of our networks. Although this ensemble network leads already to a false-positive rate (FPR) of 0.01% at a true-positive rate (TPR) of 75% on known real lenses, we have elaborated techniques to further clean the network candidate list before visual inspection. In detail, we tested the rejection using SExtractor and the modeling network from HOLISMOKES IX, which resulted together in a candidate rejection of 29% without lowering the TPR. After the initial visual inspection stage to remove obvious non-lenses, 3,408 lens candidates of the 110 million parent sample remained. We carried out a comprehensive multistage visual inspection involving eight individuals and identified finally 95 grade A (average grade ) and 503 grade B () lens candidates, including 92 discoveries showing clear lensing features that are reported for the first time. This inspection also incorporated a novel environmental characterization using histograms of photometric redshifts. We publicly release the average grades, mass model predictions, and environment characterization of all visually inspected candidates, while including references for previously discovered systems, which makes this catalog one of the largest compilation of known lenses. The results demonstrate that (1) the combination of multiple networks enhances the selection performance and (2) both automated masking tools as well as modeling networks, which can be easily applied to hundreds of thousands of network candidates expected in the near future of wide-field imaging surveys, help reduce the number of false positives, which has been the main limitation in lens searches to date.
@@
Strong gravitational lensing has emerged in recent decades as a powerful tool to probe galaxy evolution and cosmology. It allows us to obtain in a very precise way the total mass (i.e., baryonic and dark matter (DM)) of the galaxy or galaxy cluster acting as the lens (e.g., Bolton et al., 2008; Shu et al., 2017; Caminha et al., 2019) and, by assuming that mass follows light, we can disentangle the mass components and obtain unique insights into the DM distribution (e.g., Schuldt et al., 2019; Shajib et al., 2021; Wang et al., 2022) or DM substructure (e.g., Ertl et al., 2024; Lange et al., 2025; Enzi et al., 2025). Thanks to the lensing magnification, strong lensing also allows us to study high-redshift sources not visible otherwise (e.g., Shu et al., 2018; Vanzella et al., 2021; Meštrić et al., 2022; Stiavelli et al., 2023; Morishita et al., 2024).
@@
In the case of a time-variable background object, such as a supernova (SN) or a quasar, time delays can be measured (e.g., Courbin et al., 2018; Millon et al., 2020) and exploited for competitive measurements of the value of the Hubble constant, (Refsdal, 1964). Given the rarity of SNe and strong lensing events, this time-delay cosmography (TDC) technique was mostly carried out with quasars (e.g., Wong et al., 2020; Acebron et al., 2022; Shajib et al., 2022; Acebron et al., 2024). To date only three strongly lensed SNe are known with time delays usable for a precise measurement of : SN Refsdal lensed by the cluster MACS J1149.52223 (e.g., Grillo et al., 2018; Kelly et al., 2023; Grillo et al., 2024), SN H0pe strongly lensed by the cluster PLCK G165.767.0 (Frye et al., 2024; Pascale et al., 2025), and SN Encore with SN Requiem, both lensed by the same cluster, MACS J01382155 (Rodney et al., 2021; Pierel et al., 2024; Granata et al., 2025; Ertl et al., 2025, Pierel et al. in prep.). To prepare a systematic search with current and upcoming wide-field imaging surveys, we initiated the Highly Optimized Lensing Investigations of Supernovae, Microlensing Objects, and Kinematics of Ellipticals and Spirals (HOLISMOKES Suyu et al., 2020) program. As a precursor to the Legacy Survey of Space and Time (LSST) of the Vera C. Rubin Observatory (Ivezic et al., 2008; Ivezić et al., 2019), we are currently exploiting data from the Hyper Suprime-Cam (HSC), which are expected to be very similar.
@@
In Cañameras et al. (2021, hereafter C21) we presented a residual neural network to search broadly for any static lens, and complemented this in Shu et al. (2022, hereafter S22) with deflectors at relatively high redshift () to better cover the whole redshift range. Both projects relied on a single convolutional neural network (CNN) and targeted HSC images of the second public data release (PDR2). In this work, we combine multiple networks into an ensemble network, relax our restrictions on filter coverage to increase the observed sky area (requesting bands instead of ), and consider data from PDR3 with a slightly deeper and larger footprint.
@@
While the resulting sample of network candidates from HSC is small enough for visual inspection, this will not be the case for LSST, Euclid (Laureijs et al., 2011), Roman (Spergel et al., 2015), and the Chinese Space Station Telescope (Gong et al., 2019), delivering more than a billion images, even with a citizen science approach (e.g., Holloway et al., 2024; Euclid Collaboration et al., 2025a, b). Consequently, we urgently need further automated ways to lower the false-positive rate (FPR) before an unavoidable visual inspection.
@@
In this paper, we show our deep learning ensemble network and, following C21, apply it to cutouts of any object with an -Kron radius above 08 that passes standardized HSC image quality flags (see C21 for details). We explore two different approaches to reject false positives: (1) we run SExtractor to reject images with artifacts or without a real astrophysical source in the cutout, followed by a exclusion through HSC-pixel flags, and (2), we run the residual neural network from Schuldt et al. (2023a, hereafter S23a) to reject systems based on the mass model predictions. Specifically, Sect. 2 presents the tested network committee and their performances on known real lenses. We then describe the two approaches of contaminant rejection in Sect. 3, before we carry out a visual inspection as described in Sect. 4 and present the newly discovered lens candidates. Finally, we give a summary and conclusion in Sect. 5. We exploit trained networks from C21, S22, and Cañameras et al. (2024, hereafter C24), and consequently also adopt a flat CDM cosmology with = 0.32 (Planck Collaboration et al., 2020) and (Bonvin et al., 2017).
@@ 2 Inference with network committees
We followed the approach described in C21 to classify the 110 million galaxies from HSC Wide PDR3 with an -band Kron radius of 08 and optimize the contamination rate of the candidate strong lens sample. Based on test sets drawn from HSC Wide PDR2 images and designed to closely match a real classification setup, Cañameras et al. (2024) show that the purity and overall classification performance are significantly improved with committees of multiple neural networks (see also e.g., Andika et al., 2023). The highest gain is obtained when combining networks trained on different ground-truth datasets, with different prescriptions for the parameters’ distributions over the mock lenses used as positive examples. Taking the average or multiplication of output scores from networks that have little overlap in false positives due to their internal representations is best for improving the true-positive rate (TPR), defined as true-positives over all considered positives, at low FPR.
We investigated several combinations of neural networks chosen among the best performing ResNet and classical CNNs from C21, S22, and C24. All considered networks reached on their own a TPR of 40-60% at a FPR of 0.01% using 70,910 non-lens images from the COSMOS field. We refer to the original publications for more details such as their receiver operating characteristic curves.
@@
After classifying the 110 million HSC PDR3 image cutouts with these networks, we compared the TPR over three independent sets of confirmed and candidate strong lenses as a function of the total number of lens candidates predicted by the committee. This allowed us to find the optimal committee that maximizes the TPR at a given output sample size (i.e., corresponding to a fixed human inspection time). The first set of test lenses (set 1) includes the 1,249 galaxy-scale grade A or B candidates from the SuGOHI papers111see https://www-utap.phys.s.u-tokyo.ac.jp/~oguri/sugohi/. The second set (set 2) focuses on 201 galaxy-scale grade A from SuGOHI, which are also included in set 1. The third set (set 3) corresponds to the cleaned set of 178 galaxy-scale grade A or B lens candidates used in C21 C24, which excludes visually identified systems with multiple lens galaxies, or significant perturbation from the environment.
@@
We investigated various combinations of different networks by taking the average or the product of their individual network scores. As examples, the performance of combinations with two, three, four, and five networks are shown in Fig. 1. The best committee that maximizes the TPR in all three test sets at a fixed number of candidates was obtained by averaging the scores of five individual networks: the ResNet from C21, Classifier-1 from S22, and three additional ResNets from C24, as is illustrated in Fig. 2. We adopted a threshold on the average score of 0.55, which corresponds to 22,393 strong-lens candidates, and a TPR of 51%, 54%, and 75% over test set 1, 2, and 3, respectively. Since the TPR curves reach a plateau for all three of the test sets considered, decreasing the score threshold to include 50,000 additional candidates would have improved the TPR by only 4–5%, which does not justify nearly tripling the human inspection time222Using a threshold of , we obtained around 20,000 network candidates. Therefore, 50,000 additional systems would increase the system by a factor of , leading to the specified cut on the network score. Since we recorded the annotation time during our visual inspection described in Sect. 4, we also estimated – a posteriori – the actual inspection time that would have been necessary for the additional sample. Here we obtained a factor of the additional time, and possibly additional lenses.. Having various test sets was important, so that we could check that the plateau in TPR is reached irrespective of the exact setup of the test set. Since set 3 best represents the galaxy-galaxy scale lenses we target with this network committee (e.g., set 3 is cleaned from group and cluster lenses), the network committee shows the best performance on this set as expected. Finding about 20,000 candidates among 110 million sources from the parent sample is consistent with the FPR in the range of 0.01–0.03% predicted in C24 for network committees, at a TPR of 75% evaluated among set 3.
@@
The five networks are all from the same type of architecture and include residual neural network blocks (He et al., 2015). Specifically, the Classifier-1 from S22 is based on the CMU DeepLens package from Lanusse et al. (2018), while the other four networks are adapted from the ResNet-18 architecture. The networks were, however, trained on different ground-truth datasets. We introduce the general properties of the training sets and refer the reader to C21, S22, and C24 for further details.
@@
For all five networks, the realistic mock lenses used as positive examples were simulated with the pipeline described in Schuldt et al. (2021, 2023b). Briefly, the pipeline paints lensed arcs on HSC images of massive luminous red galaxies (LRGs) using singular isothermal ellipsoid (SIE) lens mass profiles, and SIE parameter values inferred from SDSS spectroscopic redshifts and velocity-dispersion measurements. After the high-redshift background sources are drawn from the Hubble Ultra-Deep Field (HUDF, Inami et al., 2017), mock lensed arcs are computed with Glee (Suyu & Halkola, 2010; Suyu et al., 2012b), and convolved with the HSC point spread function, before coaddition with the lens galaxy cutout. Positive examples used to train the ResNet from C21 and Classifier-1 from S22 were drawn from the same parent set of mocks, with (i) a nearly uniform Einstein radius distribution between 075 and 25, (ii) a boosted fraction of lens galaxies at with respect to the parent SDSS sample, (iii) lensed images that have , and (iv) a minimal ratio of between the brightest arc pixel and the local sky background over the lens LRG cutout. The other three high-performing networks from C24 that are part of the committee were trained on mocks produced with a similar procedure, but with (i) a boosted fraction of red HUDF sources, (ii) a natural Einstein radius distribution between 075 and 25 instead of an uniform distribution, and (iii) no lower limit on and a balanced fraction of double and quad configurations. This corresponds to sets labeled L1, L4, and L6 in C24.
@@
In terms of negative examples, four out of the five networks include a mix of 33% spirals, 27% isolated LRGs without arcs, 6% groups, and 33% random galaxies over the HSC footprint, as is defined for set N1 in C24. The fifth network, namely Classifier-1 from S22, was primarily targeting high-redshift strong-lenses and trained on negative examples drawn from a parent sample with red () and () colors. All five networks were trained and validated on images in bands, such that we require only the availability of bands in HSC, while the image in or can be missing.
@@ 3 Cleaning the output candidate list
Before conducting a visual inspection (see Sect. 4) to identify the high-quality strong lens candidates, the catalog of 22,393 candidates was post-processed to remove obvious artifacts and non-lenses. Since prior to this work no larger samples of HSC images were flagged as images with artifacts or crowded fields, we could not include a larger fraction of them in the negative training set. This would have helped to lower the FP rate. However, it would have lowered the fraction of other types (such as LRGs, spirals, or ring galaxies; see C24) in the negative set that more closely resemble the lens images. For this post-processing, we applied mainly two criteria, as is detailed below. An overview table of the stages and resulting sample sizes is given in Table 1. While both techniques could have also been applied to the 110 million parent sample and only the cleaned sample would have been classified by the network committee, it is significantly more efficient to first rank the full sample, and then further clean only the top candidates. This is mostly due to the longer runtime of the cleaning scripts than a network evaluation and because the second approach (see Sect. 3.2) requires in addition to the bands the band observations which can then be downloaded only for the top ranked candidates rather than the full 110 million parent sample.
@@ 3.1 Cleaning using SExtractor and HSC pixel level flags
A substantial fraction of contaminants correspond to cutouts with residual background emission in one or all bands (see some examples in the top row of Fig. 3). These cutouts were identified using source masking and estimates of the sky background with SExtractor (Bertin & Arnouts, 1996), leaving 19,820 candidates without loss of any test lens. This step was then refined with information from the pixel-level flags inferred by the HSC pipeline. We searched for optimal cuts to remove contaminants using the flags in bands over the 72 72 pixel images. First, we discarded cutouts that have 90% pixels with flags 512 and 1024, corresponding to pixels located near a bright object. This securely excluded cutouts affected by bright neighboring stars within a few tens of arcseconds. Secondly, we removed cutouts with 99% pixels flagged as “detected pixels”, which are more likely to correspond to crowded fields such as stellar clusters, or star-forming clumps within extended disks, than strong gravitational lenses. These two criteria decreased the sample to 18,712 candidates.